OpenStudy (anonymous):

Solve the systems of non linear equations.. Can you please help me walk through this, for some reason I cant seem to get how to do it still.. 2x^2-18x-y=-21 -4x^2+3x-y=-6

OpenStudy (anonymous):

for doing this sum u need to eliminate,y and then proceed to solve the quadratic equation. now corresponding to values of x u'll get value(s) for y. clue:subtract both the equations to eliminate y.

OpenStudy (anonymous):

Subtract them together?? What do you mean like use the substitution method??

OpenStudy (dumbcow):

he is referring to the elimination or addition method where you try to eliminate a variable by adding the equations (adding the coefficients)

OpenStudy (dumbcow):

you could also use substitution, solve each equation for y then set them equal to each other

OpenStudy (dumbcow):

think you get it now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I dont see how I change the soultion to y='s i need to get points for my answer (x,y) and there may be more than one set of points

OpenStudy (anonymous):

I did the substitution method and got: -2x^2+15x=-27

OpenStudy (dumbcow):

ok not quite y=2x^2 -18x +21 y=-4x^2 +3x +6 2x^2 -18x +21 = -4x^2 +3x+6 6x^2 -21x +15 = 0

OpenStudy (dumbcow):

then factor 3(2x^2 - 7x +5) =0 3(2x-5)(x-1) = 0 x=1 x=5/2 y = 2(1)^2 -18(1) +21 = 5 y=2(5/2)^2 -18(5/2) +21 = -23/2

OpenStudy (anonymous):

ok sorry been trying to work this out

OpenStudy (anonymous):

dumbcow: can u please think about my question? the sphere and light source question ?